# Beyond Bias Scores: Unmasking Vacuous Neutrality in Small Language Models

### v2

@@ Abstract
The rapid adoption of Small Language Models (SLMs) for resource constrained applications has outpaced our understanding of their ethical and fairness implications. To address this gap, we introduce the Vacuous Neutrality Framework (VaNeu), a multi-dimensional evaluation paradigm designed to assess SLM fairness prior to deployment. The framework examines model robustness across four stages - biases, utility, ambiguity handling, and positional bias over diverse social bias categories. To the best of our knowledge, this work presents the first large-scale audit of SLMs in the 0.5–5B parameter range, an overlooked “middle tier” between BERT-class encoders and flagship LLMs. We evaluate nine widely used SLMs spanning four model families under both ambiguous and disambiguated contexts. Our findings show that models demonstrating low bias in early stages often fail subsequent evaluations, revealing hidden vulnerabilities and unreliable reasoning. These results underscore the need for a more comprehensive understanding of fairness and reliability in SLMs, and position the proposed framework as a principled tool for responsible deployment in socially sensitive settings. The code is available at: https://anonymous.4open.science/r/Vacuous-Neutrality-Framework-5314/.
@@ 1 Introduction
Large Language Models (LLMs) have achieved state-of-the-art performance across a wide range of natural language processing tasks, from question answering (QA) to multilingual generation grattafiori2024llama3herdmodels; openai2024gpt4technicalreport. Trained on massive unlabelled corpora, these models excel at capturing linguistic patterns through self-supervised learning objectives such as masked language modeling devlin-etal-2019-bert. However, their scale brings two major challenges. First, LLMs are computationally expensive to deploy locally, limiting accessibility 10.1145/3604930.3605705; zhu2024surveymodelcompressionlarge. Second, their reliance on large-scale web data makes them prone to reproducing and amplifying harmful social biases, with fairness risks in high-stakes settings such as healthcare and education kaneko2021unmaskingmaskevaluating; schmidgall2024addressingcognitivebiasmedical.
To overcome the computational barrier, researchers have increasingly turned to SLMs typically under 5B parameters that offer faster inference, lower memory requirements, and reduced environmental impact. SLMs emerge either through compressing larger LLMs meta_llama_3_2_connect; gemmateam2025gemma3technicalreport, or by training compact architectures from scratch abdin2024phi3technicalreporthighly; qwen2025qwen25technicalreport. Their efficiency makes them particularly attractive for deployment on edge devices, where resources are constrained but fairness and robustness remain critical. Most SLMs rely on compression techniques such as pruning, quantization, and knowledge distillation to balance efficiency with accuracy. Yet, compression is not fairness-neutral: pruning strategies like Wanda sun2024simpleeffectivepruningapproach or SparseGPT frantar2023sparsegptmassivelanguagemodels, and quantization methods like AWQ lin2024awqactivationawareweightquantization, may inadvertently reshape model biases. This highlights the need to jointly assess performance and fairness in SLMs rather than privileging only one direction goncalves-strubell-2023-understanding.
@@
While bias and fairness evaluations have been extensively conducted on very large models (8B+) huang2023trustgptbenchmarktrustworthyresponsible; gallegos2024selfdebiasinglargelanguagemodels and smaller models under 0.5B parameters such as BERT parrish-etal-2022-bbq, the intermediate range of 0.5B–5B remains largely understudied-despite its growing significance for practical deployment. These mid-sized models strike a balance between efficiency and capability, making them especially relevant for real-world applications. This gap raises an important question: Can these SLMs be trusted in socially sensitive settings?
@@ 2 Related Work
Numerous studies have shown that LLMs not only reflect existing social biases in their responses, particularly around sensitive attributes such as gender, race, and sexual orientation but can also amplify these biases during downstream tasks venkit2023nationalitybiastextgeneration; goncalves-strubell-2023-understanding. To evaluate such risks, several benchmarks have been developed, including StereoSet nadeem2020stereosetmeasuringstereotypicalbias and UNQOVER li-etal-2020-unqovering. Analyses of prominent transformer-based models such as BERT devlin2019bertpretrainingdeepbidirectional, RoBERTa liu2019robertarobustlyoptimizedbert, GPT-2 Radford2019LanguageMA, and GPT-4 törnberg2023chatgpt4outperformsexpertscrowd reveal that, despite architectural advancements and mitigation strategies such as fine-tuning or data filtering, notable biases persist. These findings highlight that fairness challenges remain deeply embedded across model families and scales.
@@
Model compression techniques, while essential for improving efficiency, can have unintended consequences for fairness. Some studies show that compression strategies exacerbate social biases in language models ramesh-etal-2023-comparative and cause unpredictable shifts in behavior xu2024perplexitymultidimensionalsafetyevaluation, whereas others suggest compression may act as a regularizer, mitigating bias in certain contexts Lin_2024. This duality arises because compression can either reduce overfitting and thereby dampen bias, or distort learned representations in ways that amplify it. Thus, the fairness implications of compression are complex and highly context-dependent.
While numerous studies confirm the persistence of social bias in LLMs gallegos2024biasfairnesslargelanguage; Li2023ASO, relatively little is known about how these biases manifest in SLMs. Existing work has predominantly focused on large-scale models (8B+ parameters) hong2024decodingcompressedtrustscrutinizing or on much smaller models such as BERT (under 0.5B parameters) goncalves-strubell-2023-understanding. This leaves a significant gap in understanding mid-sized SLMs (0.5B–5B), a model class that is increasingly attractive for deployment due to its balance of efficiency and capability. To address this gap, we conduct a systematic evaluation of open-source, transformer-based SLMs within this intermediate range, focusing specifically on their tendencies to exhibit social bias under socially sensitive benchmarks. To the best of our knowledge, this is the first comparative fairness audit spanning multiple transformer families of SLMs in the 0.5B–5B parameter range across widely used bias evaluation benchmarks.
3 Vacuous Neutrality Framework
Evaluating SLMs requires going beyond single dimension metrics. We introduce the Vacuous Neutrality Framework (VaNeu), a multi-stage evaluation paradigm designed to assess SLMs across four complementary dimensions: Bias, Utility, Ambiguity Handling, and Positional Bias.
@@ 4.2 Datasets
BBQ (Bias Benchmark for QA) parrish-etal-2022-bbq: A large-scale question answering dataset designed to test stereotypical reasoning under both ambiguous and disambiguated contexts. Each instance pairs a question with demographic attributes such as gender, race, religion, or nationality. Ground truth labels are provided at the question level, which enables direct evaluation of both bias (e.g., Bias Score) and utility (e.g., Accuracy and F1 Score). BBQ is also the only dataset among the three that natively supports ambiguity handling, since it includes cases where the correct answer is “Unknown.”
StereoSet nadeem2020stereosetmeasuringstereotypicalbias: A benchmark for measuring stereotypical bias in natural language understanding. Each context is paired with candidate completions that may be stereotypical, anti-stereotypical, or unrelated. Ground truth is provided only at the level of stereotypicality, that is, whether a completion reflects a stereotype, an anti-stereotype, or an unrelated association, rather than specifying a task-correct answer. This structure makes StereoSet well-suited for evaluating bias tendencies, but less informative for measuring utility or ambiguity handling without modification.
CrowS-Pairs nangia2020crows: A minimal-pair dataset where each instance contrasts a biased and an unbiased alternative differing only by a single lexical substitution. Ground truth is provided only at the level of stereotype polarity, whether a sentence is stereo or anti-stereo, rather than specifying a task-correct answer. This design enables precise bias quantification, but does not natively support evaluation of utility or ambiguity handling.
@@ 4.3 Evaluation Metrics
Utility Dimension For StereoSet and CrowS-Pairs, we evaluate utility using the Language Modeling Score (LMS) nadeem2020stereosetmeasuringstereotypicalbias, defined as the percentage of instances where the model favors a meaningful (stereotypical or anti-stereotypical) association over an unrelated one. An ideal model attains an LMS of 100. For BBQ, we measure task performance using the F1 score, computed separately for ambiguous and disambiguated contexts.
@@
where $P(i)$ is the predicted probability for position $i$, $Q(i)$ is the ground truth or uniform distribution, and $|C|$ is the number of classes.
@@ 5 Experiments and Results
By contrast, the Phi family demonstrates strong robustness. Phi-4-Mini maintains a balanced target-to-nontarget ratio across most categories (except minor deviations in Disability Status and Religion) and aligns closely with the ground-truth unknown distribution, except for Age and Disability Status. This reflects an ability to abstain when necessary without compromising task competence. Phi-3.5-Mini exhibits similar and even stronger stability, though with slightly greater variability across categories. Based on Stage 3, Phi-3.5-Mini and Phi-4-Mini maintain balanced caution and specificity, advancing to the final stage.
@@ 6 Discussion
If deployment were based only on Stages 1 and 2, we would risk releasing biased or unstable models. As discussed in Appendix A.2, Qwen2.5-3B initially appears deployable after Stage 1 but exhibits an extremely high TNR (153.86) in Disability Status Category and consistently low Norm-DKL (< 0.10) across categories, indicating overcommitment to a single option and a lack of meaningful differentiation. LLaMA3.2-3B performs in the utility under disambiguated contexts but fails in Stage 3, showing poor UR calibration and strong positional preference in Stage 4. Similarly, Gemma3-4B achieves high task utility in disambuigated contexts yet struggles with ambiguity handling. However, its Stage 4 answer distribution aligns more closely with the ground truth, suggesting that apparent neutrality stems from balanced outputs rather than genuine reasoning. We further tested the effect of task-specific fine-tuning (Appendix D); it improved disambiguated performance but reduced reasoning under ambiguity. Stages 3 and 4 refine the analysis by assessing specificity and distributional balance, revealing that fairness and utility must be interpreted jointly, as models prone to vacuous neutrality may appear reliable without genuine reasoning. We discussed the stage 3 and stage 4 results for Small LMs (2B-4B) in the Appendix  A.2
@@ 7 Conclusion
8 Limitations
Our study is subject to several limitations that warrant consideration and highlight avenues for future research. First, we focus exclusively on open-source SLMs within the 0.5B–5B parameter range. Consequently, our observations on bias–capacity trade-offs are limited to this intermediate scale and may not extend to larger or proprietary models such as GPT-4 (OpenAI et al., 2024). Second, our evaluation relies on the BBQ dataset (Parrish et al., 2022), which effectively probes bias under contextual ambiguity but remains confined to U.S.-centric social categories and a question-answering format. Extending the framework to multicultural contexts, different languages, additional architectures and broader downstream tasks such as summarization, dialogue, or retrieval would enhance generalizability. Finally, the notion of Vacuous Neutrality remains qualitatively defined; future work aims to formalize it mathematically and integrate the four stages into a unified composite metric.
9 Ethical Considerations
@@ Appendix A BBQ Dataset
The Bias Benchmark for Question Answering (BBQ) dataset parrish-etal-2022-bbq is a comprehensive benchmark designed to assess representational biases in language models. The BBQ dataset is licensed for non-commercial research use. All evaluated models are publicly available under open-source licenses (e.g., Apache 2.0, MIT) via HuggingFace. It comprises 58,492 unique question instances, each presented in both ambiguous and disambiguated formats. The dataset covers nine key demographic dimensions and two intersectional dimensions to facilitate a deeper examination of compound biases. Each question presents three answer choices: one that reflects a stereotypical bias (Target), one that challenges the stereotype (Non-Target), and an “Unknown” choice that reflects appropriate uncertainty. To evaluate model behavior, the original authors propose four metrics: accuracy on ambiguous questions (where the correct response is ideally “Unknown”), accuracy on disambiguated questions (where the model is expected to select the contextually appropriate answer), and two bias scores quantifying stereotypical tendencies under both ambiguous, $s_{\text{AMB}}$ and disambiguated conditions, $s_{\text{DIS}}$. In this paper, we adopt the F1 score in place of accuracy to evaluate the utility of the model. Both bias scores falls within the range [100,+100][-100,+100], where values near zero indicate low bias or neutral.

### v3

@@ Abstract
The rapid adoption of Small Language Models (SLMs) for resource constrained applications has outpaced our understanding of their ethical and fairness implications. To address this gap, we introduce the Vacuous Neutrality Framework (VaNeu), a multi-dimensional evaluation paradigm designed to assess SLM fairness prior to deployment. The framework examines model robustness across four stages - biases, utility, ambiguity handling, and positional bias over diverse social bias categories. To the best of our knowledge, this work presents the first large-scale audit of SLMs in the 0.5–5B parameter range, an overlooked “middle tier” between BERT-class encoders and flagship LLMs. We evaluate nine widely used SLMs spanning four model families under both ambiguous and disambiguated contexts. Our findings show that models demonstrating low bias in early stages often fail subsequent evaluations, revealing hidden vulnerabilities and unreliable reasoning. These results underscore the need for a more comprehensive understanding of fairness and reliability in SLMs, and position the proposed framework as a principled tool for responsible deployment in socially sensitive settings. The code is available at: https://github.com/smanduru10/Vacuous-Neutrality-Framework.git.
@@ 1 Introduction
Large Language Models (LLMs) have achieved state-of-the-art performance across a wide range of natural language processing tasks, from question answering (QA) to multilingual generation (Grattafiori et al., 2024; OpenAI et al., 2024). Trained on massive unlabelled corpora, these models excel at capturing linguistic patterns through self-supervised learning objectives such as masked language modeling (Devlin et al., 2019a). However, their scale brings two major challenges. First, LLMs are computationally expensive to deploy locally, limiting accessibility (Chien et al., 2023; Zhu et al., 2024). Second, their reliance on large-scale web data makes them prone to reproducing and amplifying harmful social biases, with fairness risks in high-stakes settings such as healthcare and education (Kaneko and Bollegala, 2021; Schmidgall et al., 2024).
To overcome the computational barrier, researchers have increasingly turned to SLMs typically under 5B parameters that offer faster inference, lower memory requirements, and reduced environmental impact. SLMs emerge either through compressing larger LLMs (Llama3.2, 2024; GemmaTeam et al., 2025), or by training compact architectures from scratch (Abdin et al., 2024; Qwen et al., 2025). Their efficiency makes them particularly attractive for deployment on edge devices, where resources are constrained but fairness and robustness remain critical. Most SLMs rely on compression techniques such as pruning, quantization, and knowledge distillation to balance efficiency with accuracy. Yet, compression is not fairness-neutral: pruning strategies like Wanda (Sun et al., 2024) or SparseGPT (Frantar and Alistarh, 2023), and quantization methods like AWQ (Lin et al., 2024a), may inadvertently reshape model biases. This highlights the need to jointly assess performance and fairness in SLMs rather than privileging only one direction (Gonçalves and Strubell, 2023).
@@
While bias and fairness evaluations have been extensively conducted on very large models (8B+) (Huang et al., 2023; Gallegos et al., 2024b) and smaller models under 0.5B parameters such as BERT (Parrish et al., 2022), the intermediate range of 0.5B–5B remains largely understudied-despite its growing significance for practical deployment. These mid-sized models strike a balance between efficiency and capability, making them especially relevant for real-world applications. This gap raises an important question: Can these SLMs be trusted in socially sensitive settings?
@@ 2 Related Work
Numerous studies have shown that LLMs not only reflect existing social biases in their responses, particularly around sensitive attributes such as gender, race, and sexual orientation but can also amplify these biases during downstream tasks (Venkit et al., 2023; Gonçalves and Strubell, 2023). To evaluate such risks, several benchmarks have been developed, including StereoSet (Nadeem et al., 2020) and UNQOVER (Li et al., 2020). Analyses of prominent transformer-based models such as BERT (Devlin et al., 2019b), RoBERTa (Liu et al., 2019), GPT-2 (Radford et al., 2019), and GPT-4 (törnberg2023chatgpt4outperformsexpertscrowd) reveal that, despite architectural advancements and mitigation strategies such as fine-tuning or data filtering, notable biases persist. These findings highlight that fairness challenges remain deeply embedded across model families and scales.
@@
Model compression techniques, while essential for improving efficiency, can have unintended consequences for fairness. Some studies show that compression strategies exacerbate social biases in language models (Ramesh et al., 2023) and cause unpredictable shifts in behavior (Xu et al., 2024), whereas others suggest compression may act as a regularizer, mitigating bias in certain contexts (Lin et al., 2024b). This duality arises because compression can either reduce overfitting and thereby dampen bias, or distort learned representations in ways that amplify it. Thus, the fairness implications of compression are complex and highly context-dependent.
While numerous studies confirm the persistence of social bias in LLMs (Gallegos et al., 2024a; Li et al., 2023), relatively little is known about how these biases manifest in SLMs. Existing work has predominantly focused on large-scale models (8B+ parameters) (Hong et al., 2024) or on much smaller models such as BERT (under 0.5B parameters) (Gonçalves and Strubell, 2023). This leaves a significant gap in understanding mid-sized SLMs (0.5B–5B), a model class that is increasingly attractive for deployment due to its balance of efficiency and capability. To address this gap, we conduct a systematic evaluation of open-source, transformer-based SLMs within this intermediate range, focusing specifically on their tendencies to exhibit social bias under socially sensitive benchmarks. To the best of our knowledge, this is the first comparative fairness audit spanning multiple transformer families of SLMs in the 0.5B–5B parameter range across widely used bias evaluation benchmarks.
3 The Vacuous Neutrality Framework
Evaluating SLMs requires going beyond single dimension metrics. We introduce the Vacuous Neutrality Framework (VaNeu), as shown in Figure 1, a multi-stage evaluation paradigm designed to assess SLMs across 4 complementary dimensions: Bias, Utility, Ambiguity Handling, and Positional Bias.
Vacuous Neutrality:
We define vacuous neutrality as a failure mode in which a language model attains low measured bias under bias-centric evaluation while lacking the competence, calibration, or robustness required for reliable reasoning. Formally, a model exhibits vacuous neutrality when apparent neutrality arises not from principled inference, but from degenerate behaviors such as random guessing, indiscriminate abstention, over-commitment to a single option, or reliance on superficial heuristics. In such cases, low bias scores coexist with poor task utility, uncalibrated uncertainty under ambiguity, or artifact-driven decision patterns, rendering the model unreliable for deployment despite its ostensibly fair behavior.
@@ 4.2 Datasets
BBQ (Bias Benchmark for QA) (Parrish et al., 2022): A large-scale QA dataset designed to test stereotypical reasoning under both ambiguous and disambiguated contexts. Each instance pairs a question with demographic attributes such as gender, race, religion, or nationality. Ground truth labels are provided at the question level, which enables direct evaluation of both bias (e.g., Bias Score) and utility (e.g., Accuracy and F1 Score). BBQ is also the only dataset among the three that natively supports ambiguity handling, since it includes cases where the correct answer is “Unknown.”
StereoSet (Nadeem et al., 2020): A benchmark for measuring stereotypical bias in natural language understanding. Each context is paired with candidate completions that may be stereotypical, anti-stereotypical, or unrelated. Ground truth is provided only at the level of stereotypicality, that is, whether a completion reflects a stereotype, an anti-stereotype, or an unrelated association, rather than specifying a task-correct answer. This structure makes StereoSet well-suited for evaluating bias tendencies, but less informative for measuring utility or ambiguity handling without modification.
CrowS-Pairs (Nangia et al., 2020): A minimal-pair dataset where each instance contrasts a biased and an unbiased alternative differing only by a single lexical substitution. Ground truth is provided only at the level of stereotype polarity, whether a sentence is stereo or anti-stereo, rather than specifying a task-correct answer. This design enables precise bias quantification, but does not natively support evaluation of utility or ambiguity handling.
@@ 4.3 Evaluation Metrics
Utility Dimension For StereoSet and CrowS-Pairs, we evaluate utility using the Language Modeling Score (LMS) (Nadeem et al., 2020), defined as the percentage of instances where the model favors a meaningful (stereotypical or anti-stereotypical) association over an unrelated one. An ideal model attains an LMS of 100. For BBQ, we measure task performance using the F1 score, computed separately for ambiguous and disambiguated contexts.
@@
where $P(i)$ is the predicted probability for position $i$, $Q(i)$ is the ground truth or uniform distribution, and $|C|$ is the number of classes. Refer to Appendix E for additional discussion.
@@ 5 Experiments and Results
By contrast, the Phi family demonstrates strong robustness. Phi-4-Mini maintains a balanced target-to-nontarget ratio across most categories (except minor deviations in Disability Status and Religion) and aligns closely with the ground-truth unknown distribution, except for Age and Disability Status. This reflects an ability to abstain when necessary without compromising task competence. Phi-3.5-Mini exhibits similar and even stronger stability, though with slightly greater variability across categories. Based on Stage 3, both Phi Models maintain balanced caution and specificity, advancing to the final stage.
@@ 6 Discussion
If deployment were based only on Stages 1 and 2, we would risk releasing biased or unstable models. As discussed in Appendix A.2, Qwen2.5-3B initially appears deployable after Stage 1 but exhibits an extremely high TNR (153.86) in Disability Status Category and consistently low Norm-DKL (< 0.10) across categories, indicating overcommitment to a single option and a lack of meaningful differentiation. LLaMA3.2-3B performs in the utility under disambiguated contexts but fails in Stage 3, showing poor UR calibration and strong positional preference in Stage 4. Similarly, Gemma3-4B achieves high task utility in disambuigated contexts yet struggles with ambiguity handling. However, its Stage 4 answer distribution aligns more closely with the ground truth, suggesting that apparent neutrality stems from balanced outputs rather than genuine reasoning. We further tested the effect of task-specific fine-tuning (Appendix D); it improved disambiguated performance but reduced reasoning under ambiguity. Stages 3 and 4 refine the analysis by assessing specificity and distributional balance, revealing that fairness and utility must be interpreted jointly, as models prone to vacuous neutrality may appear reliable without genuine reasoning. We discussed the results of stages 3 and 4 for SLMs (2B-4B) in the Appendix  A.2
@@
Bias-Centric Benchmarks under VaNeu:
To contextualize how bias-centric audits relate to the VaNeu Framework, we evaluate four SLMs: LLaMA-3.2-3B, Gemma3-4B, Phi-3.5-mini, and Phi-4-mini on StereoSet and CrowS-Pairs (Appendix B, Appendix C). Under standard reporting on these benchmarks, all four models appear broadly acceptable. Stereo Scores are generally moderate, Language Modeling Scores are often high, and the S/AS/U distributions indicate that models typically produce non-unrelated completions with some degree of abstention. However, because StereoSet and CrowS-Pairs provide supervision primarily for directional social bias (stereotypical versus anti-stereotypical preference) and do not supply task-correct answers, explicit ambiguity control, or reference distributions for positional robustness, these results are necessary but insufficient for deployment decisions. In particular, such metrics cannot distinguish principled neutrality from conservative or heuristic behavior (e.g., over-commitment, elevated Unknown usage, or superficially balanced outputs that still score well on SS/LMS/iCAT). This limitation motivates VaNeu’s staged design, when the same models are assessed using a benchmark that supports competence and ambiguity evaluation (i.e., BBQ), models that appear similarly well-behaved under bias-only metrics separate sharply in reliability, revealing brittleness or inefficiency for some (e.g., LLaMA-3.2-3B and Gemma3-4B) and more robust behavior for others (Phi-4-mini, with Phi-3.5-mini exhibiting intermediate robustness). More broadly, these findings suggest that existing bias benchmarks are insufficient to diagnose vacuous neutrality in isolation. Extending VaNeu beyond BBQ will therefore require complementary datasets that explicitly control ambiguity, provide per-instance ground truth, and balance answer positions, enabling joint evaluation of bias, utility, ambiguity handling, and positional robustness in socially sensitive settings.
@@ 7 Conclusion
Limitations
Our study is subject to several limitations that warrant consideration and highlight avenues for future research. First, we focus exclusively on open-source SLMs within the 0.5B–5B parameter range. Consequently, our observations on bias–capacity trade-offs are limited to this intermediate scale and may not extend to larger or proprietary models such as GPT-4 (OpenAI et al., 2024). Second, our evaluation is conducted on bias-related datasets designed to probe contextual ambiguity, but these datasets are largely limited to U.S.-centric social categories and a question-answering format. Extending the framework to multilingual and multicultural settings, alternative architectures, and broader downstream tasks such as summarization, dialogue, or retrieval would further enhance its generalizability. Finally, while Vacuous Neutrality is operationalized through a set of quantitative stages, an important direction for future work is to formalize this notion mathematically and integrate the stages into a unified composite metric.
Ethical Considerations
@@ Appendix A BBQ Dataset
The Bias Benchmark for Question Answering (BBQ) dataset (Parrish et al., 2022) is a comprehensive benchmark designed to assess representational biases in language models. The BBQ dataset is licensed for non-commercial research use. All evaluated models are publicly available under open-source licenses (e.g., Apache 2.0, MIT) via HuggingFace. It comprises 58,492 unique question instances, each presented in both ambiguous and disambiguated formats. The dataset covers nine key demographic dimensions and two intersectional dimensions to facilitate a deeper examination of compound biases. Each question presents three answer choices: one that reflects a stereotypical bias (Target), one that challenges the stereotype (Non-Target), and an “Unknown” choice that reflects appropriate uncertainty. To evaluate model behavior, the original authors propose four metrics: accuracy on ambiguous questions (where the correct response is ideally “Unknown”), accuracy on disambiguated questions (where the model is expected to select the contextually appropriate answer), and two bias scores quantifying stereotypical tendencies under both ambiguous, $s_{\text{AMB}}$ and disambiguated conditions, $s_{\text{DIS}}$. In this paper, we adopt the F1 score in place of accuracy to evaluate the utility of the model. Both bias scores falls within the range [100,+100][-100,+100], where values near zero indicate low bias or neutral.